The temperature coefficient of resistivity of a material is 0.0036 (C°)-1. By how much ΔT does the temperature of a wire made from this material have to change, in order that the resistance of the wire doubles?
Answer #1

α = 0.0036 / C
Temperature dependence of Resistance of a material is given by,
R = Rref * [ 1 + α (ΔT) ]
For Resistance to double,
2R = R * [ 1 + α (ΔT) ]
2 = 1 + α *ΔT
ΔT = 1 / 0.0036
ΔT = 277.8 oC